OpenStudy (eric_d):

Example: i)(x+a)(x+b)=x^2+x(a+b)+ab ii) (a+b)^2=... There are many more Can you help to list it down...

OpenStudy (eric_d):

@ikram002p

OpenStudy (ikram002p):

ok ? what is ur Qn exactly ?

OpenStudy (eric_d):

I just want to know all the possible formula* For instance (a+b)^2= a+2ab+c

ganeshie8 (ganeshie8):

you don't need to memorize the first formula

ganeshie8 (ganeshie8):

(x+a)(x+b) FOIL this.

OpenStudy (ikram002p):

\(( a+b)^2=a^2+2ab+b^2 \) ?

OpenStudy (ikram002p):

OH my !

ganeshie8 (ganeshie8):

i memorize only below formula : 1) \(\large (a+b)^2 = a^2 + 2ab + b^2\) 2) \(\large (a-b)^2 = a^2 - 2ab + b^2\) 3) \(\large (a+b+c)^2 = a^2 +b^2+c^2+2ab+2bc+2ca\)

ganeshie8 (ganeshie8):

Oh, and one more very important identity : 4) \(\large a^2-b^2 = (a+b)(a-b)\)

OpenStudy (ikram002p):

ok so his Qn what what formula he need to memorize ?

OpenStudy (eric_d):

Yes @ikram002p

OpenStudy (ikram002p):

ic ! i only memorize binomial thm :P jk only earn FOIL concept , do u know it ?

OpenStudy (ikram002p):